A sweater cost a dealer $32.00. At what price should he mark the sweater so he can give a discount of 25% and still make a mark-up profit of 20%? What is the final price if there is a 5% sales tax?
Answer by rfer(16322)

32*1.2=$38.40 sale price
38.40*1.25=48.00 full price.
38.40*1.05=$40.32 sale price with tax included.